Suppose you have two capacitors 8.0 μF and 4.0 μF connected in series and parallel with points a and b whose potential


difference is 20 V. Find the equivalent capacitance and charge.

1
Expert's answer
2022-03-01T11:12:25-0500

Parallel connection:

The equivalent capacitance

"C=C_1+C_2\\
=\rm 8.0\: \mu F+\rm 4.0\: \mu F=\rm 12\: \mu F"

The equivalent charge

"q=CV\\
=12*10^{-6}*20=240*10^{-6}\: \rm C"

Series connection:

The equivalent capacitance

"C=C_1C_2/(C_1C_2)\\
=\rm 8.0\: \mu F*\rm 4.0\: \mu F/(\rm 8.0\: \mu F+\rm 4.0\: \mu F)=\rm 2.7\: \mu F"

The equivalent charge

"q=CV\\
=2.7*10^{-6}*20=54*10^{-6}\: \rm C"